1. You purchased 280 shares of stock at a price of $53.37 per share. Over the...

1.

You purchased 280 shares of stock at a price of $53.37 per share. Over the last year, you have received total dividend income of $305. What is the dividend yield?

5.7 percent

6.2 percent

1.1 percent

16.0 percent

2.0 percent

2.

You own a stock portfolio invested 25 percent in Stock Q, 25 percent in Stock R, 15 percent in Stock S, and 35 percent in Stock T. The betas for these four stocks are 0.96, 1.24, 0.72, and 1.68, respectively. What is the portfolio beta?


rev: 09_20_2012

1.27

1.18

1.31

1.22

1.25
Answer #1

1.Total beginning price=(280*53.37)=$14943.6

Hence dividend yield=Dividend/beginning price

=(305/14943.6)

=2.0%(Approx).

2.Portfolio beta=Respective beta*Respective investment weight

=(0.25*0.96)+(0.25*1.24)+(0.15*0.72)+(0.35*1.68)

=1.25(Approx).